OpenStudy (anonymous):

A store had 250 bottles of water. Each week, 40% of the bottles were sold and 48 new bottles arrived in shipments. Which recursive function best represents the number of bottles of water in the store, given that f(0) = 250?

OpenStudy (anonymous):

So the zeroth week in the recursion would be 250 bottles, what happens after that?

OpenStudy (anonymous):

well each week 48 new bottles would come in and 40% would be taken away. These are the answers to choose from: f(n) = f(n - 1) • 0.6 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.4 + 48, n > 0 f(n) = f(n - 1) • 0.4 + 48, n > 0 f(n) = 250 - f(n - 1) • 0.6 + 48, n > 0

OpenStudy (anonymous):

I feel that I understand it but at the same time I don't...? If that makes sense

OpenStudy (anonymous):

I feel the same way haha. I am trying to wrap my head around what the right answer is before I try to walk you through it.

OpenStudy (anonymous):

Would it cancel out B and D?

OpenStudy (anonymous):

Well there are two says to look at this You have f(0) = 250 and f(1) can either equal to 250-250*40+48 In this form you have your initial and are getting rid of your 40 % sold and adding 48 or f(1) = 250 * .6 + 48 In this form you are keeping the 60 % not sold and adding 48 So my previous statement was wrong about getting rid of two (sorry).

OpenStudy (anonymous):

So because you want this to go on for more than one week yes it would get rid of B and D (sorry again).

OpenStudy (anonymous):

Since its going by 40% would it have to be C?

OpenStudy (anonymous):

Yes it would be.

OpenStudy (anonymous):

Okay thank you so much!

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

C was not the correct answer